III. Suppose that after making 50 payments, the interest rate changes to J2=9% p.a.:

A. Convert the interest rate J2=9% to J12 equivalent
1
Expert's answer
2020-06-16T19:24:27-0400
"J1=(1+{j2\over2})^2-1=(1+0.045)^2"


"=0.092025"



"J12=12((1+{0.092025\over1})^{1\over12}-1)"

"i=0.088357=8.8357\%"